# MoveGPT: Scaling Mobility Foundation Models with Spatially-Aware Mixture of Experts

### v2

@@ 1 Introduction
Understanding human mobility patterns is a cornerstone for gaining insight into complex human behaviors and socio-economic dynamics (Song et al., 2010a; Gonzalez et al., 2008). The vast digital traces from billions of personal devices have provided researchers with massive datasets of human mobility, capturing daily activities on an unprecedented scale (Feng et al., 2018; Yuan et al., 2022; Xue et al., 2021). The effort to find predictable patterns within this apparent randomness is not new; nearly two decades ago, studies revealed surprising regularities and high predictability in human mobility (Song et al., 2010b; Schneider et al., 2013). This combination of massive data and intrinsic predictability creates the ideal conditions to move beyond task-specific statistical models and towards a more holistic and generalizable understanding of human movement (Zhou et al., 2024).
Inspired by the transformative success of foundation models in language, a new wave of general-purpose models for mobility has emerged, such as UniST (Yuan et al., 2024a), UrbanDiT (Yuan et al., 2024b), TrajFM (Lin et al., 2024), TrajGPT (Hsu et al., 2024), UniMob (Long et al., 2025b), GenMove (Long et al., 2025a), and UniTraj (Zhu et al., 2024). However, these pioneering efforts, while valuable, fall short of creating a true “GPT for mobility”. A detailed comparison, summarized in Table 1, reveals several practical limitations shared by current approaches. Specifically, they rely on inconsistent basic units to represent movement, from raw GPS coordinates to non-transferable Zone IDs, a fragmentation that directly hinders their geographic scalability and prevents cross-city pre-training. Furthermore, many of these models exhibit limited task flexibility and are constrained in their parameter scale, which prevents them from fully leveraging the rich patterns within large, diverse mobility datasets.
@@
TrajFM (Lin et al., 2024) Scratch Raw GPS Irregular 18M
@@
TrajGPT (Hsu et al., 2024) Scratch Stay Point Irregular 20w
TrajGDM (Chu et al., 2023) Scratch Stay Point Irregular 28M
UniMove (Han et al., 2025) Scratch Stay Point Irregular 40M
GenMove (Long et al., 2025a) Scratch Zone ID Regular 21M
UniMob (Long et al., 2025b) Scratch Zone ID Regular 6M
UniST (Yuan et al., 2024a) Scratch Zone ID Regular 30M
UrbanGPT (Li et al., 2024b) LLMs Zone ID Regular -
UrbanDiT (Yuan et al., 2024b) Scratch Zone ID Regular 30M
@@
We argue these practical limitations stem from two fundamental yet unsolved barriers. The first barrier is the lack of a universal basic unit, which directly prevents the scaling of data across geographies. For instance, some models define tokens as aggregated spatio-temporal patches (e.g., UniST (Yuan et al., 2024a), UrbanDiT (Yuan et al., 2024b)), an approach too coarse to capture the semantics of individual movement. Others use raw GPS coordinates (e.g., UniTraj (Zhu et al., 2024)), which are too fragmented to represent daily routines, or rely on city-specific Zone IDs (e.g., TrajGPT (Hsu et al., 2024), UniMob (Long et al., 2024), GenMove (Long et al., 2025a)), which are not universal and prevent the creation of large-scale, cross-city pre-training datasets. The second barrier is architectural. Human mobility, much like language, is a mixture of distinct behaviors shaped by varied intents and geographical contexts (Han et al., 2025). Existing models, with their ”one-fits-all” parameters, are forced to learn an inefficient, averaged representation of these diverse patterns, failing to capture the rich semantics of movement as data diversity grows. Recent LLM-based approaches (Wang et al., 2023; Li et al., 2024a; Feng et al., 2025; Du et al., 2024; 2025) sidestep this by reasoning in a language space, a process that divorces the model from the data’s intrinsic spatio-temporal structure.
@@ 3.1 Unified location encoder
Existing approaches typically rely on discrete IDs to encode locations. However, such representations hinder cross-city transferability. To build a mobility foundation model, it is essential to develop a unified encoder that can generate consistent representations of locations across different cities. Existing research (Hu et al., 2020; Song et al., 2010c; Gonzalez et al., 2008; Lenormand et al., 2020) shows evidence that human mobility is influenced by individual intents, travelers tend to choose their next destination based on travel needs and spatial distance. To capture this behavior, we use the locations’ functional feature (POI distribution) and geographic feature (geographic coordinates) as core features for the encoder. In addition, mobility patterns are also influenced by broader social dynamics. To account for this, we incorporate the social popularity level of each location as another key feature. Formally, for all locations in a city, we construct three embeddings:
@@ 3.3 Next-step Output
Instead of direct classification, we model next-step prediction as a similarity search problem. Specifically, a mobility trajectory $S$ is encoded by the SAMoE Transformer to produce a trajectory representation $H_{\text{traj}}^{l}\in\mathbb{R}^{d}$ for a given length $l$, which acts as a query. Concurrently, we create a candidate database $\mathcal{D}\in\mathbb{R}^{N\times d}$ containing embeddings for all $N$ locations in the city, generated by a Deep & Cross Network (DCN) (Wang et al., 2017) from their features ($E_{\text{poi}}$, $E_{\text{geo}}$, $E_{\text{pop}}$) as follows:
@@ B.1 Human Mobility Modeling
Human mobility modeling encompasses a broad spectrum of tasks, including mobility prediction, generation, and anomaly detection. These tasks jointly provide complementary perspectives for understanding, simulating, and evaluating human movement behaviors (Song et al., 2010b; Schneider et al., 2013; Fang et al., 2025). Early approaches were dominated by probabilistic models such as Markov chains and exploration-preferential return (EPR) models (Gambs et al., 2012; Jiang et al., 2016), which captured mobility patterns from transition probabilities but struggled with the irregular and dynamic nature of human behavior. Recent advances in deep learning have significantly boosted the field, with recurrent neural networks (RNNs) (Chen et al., 2020; Yuan et al., 2022), attention-based architectures (Feng et al., 2018; Yuan et al., 2023), Transformers (Corrias et al., 2023; Si et al., 2023), graph neural networks (GNNs) (Sun et al., 2021; Terroso-Sáenz & Muñoz, 2022), and diffusion models (Zhu et al., 2023; Long et al., 2025a) being widely applied across different tasks. However, most of these models remain city-specific, requiring separate training for each dataset and thus limiting their generalizability. Inspired by the paradigm shift in natural language processing, a series of mobility foundation models have recently been proposed, including UniST (Yuan et al., 2024a), UrbanDiT (Yuan et al., 2024b), TrajFM (Lin et al., 2024), TrajGPT (Hsu et al., 2024), UniMob (Long et al., 2025b), GenMove (Long et al., 2025a), and UniTraj (Zhu et al., 2024). These approaches attempt to leverage large-scale pre-training to improve few-shot and zero-shot performance, yet most still can not capture the universality and diversity of human mobility patterns across urban environments.
@@ B.2 Mixture of Experts
Mixture of Experts (MoE)  (Shazeer et al., 2017; Jordan & Jacobs, 1994; Jacobs et al., 1991) is a deep learning architecture that enhances model performance through a divide-and-conquer strategy. Its core idea lies in decomposing complex tasks and enabling collaborative processing by multiple specialized sub-models (experts), guided by a learnable gating network that dynamically allocates input data to the most relevant experts (Zhou et al., 2022). Each expert processes only a subset of inputs, significantly reducing computational costs, while the gating mechanism synthesizes final results through weighted aggregation of expert outputs (Fedus et al., 2022; Lepikhin et al., ). By maintaining the model’s parameter scale while improving inference efficiency, MoE has been widely adopted in large-scale language models (Liu et al., 2024; Du et al., 2022; Lepikhin et al., ), enabling trillion-parameter model deployment via sparse activation. However, challenges such as expert load balancing and training stability persist in its implementation.
@@ D.1 Baselines
  • Markov (Gambs et al., 2012): The Markov model provides a statistical framework for characterizing how states evolve over time. It predicts future locations by estimating the transition probabilities between states.
  • LSTM (Graves, 2012): Long Short-Term Memory networks are highly effective for sequential modeling, as they capture long-term dependencies and temporal dynamics, making them particularly suitable for location prediction tasks.
  • Transformer (Vaswani, 2017): The Transformer architecture leverages self-attention mechanisms to process sequential data efficiently, enabling the modeling of long-range dependencies in human mobility patterns.
  • DeepMove (Feng et al., 2018): DeepMove is designed to address the challenges of human mobility prediction by jointly considering spatial, temporal, and personal preference factors that shape individual movement behaviors.
  • GetNext (Yang et al., 2022): GETNext integrates Graph Convolution Networks with Transformers to model global POI transition patterns, incorporating time-aware semantic embeddings for next-location recommendation.
  • CTLE (Lin et al., 2021): The Context- and Time-aware Location Embedding model dynamically generates location embeddings by combining bidirectional Transformers with temporal encoding, capturing multi-faceted semantics of locations under varying spatial-temporal dynamics.
  • Unimove (Han et al., 2025): Unimove is a unified model for multi-city human mobility prediction, employing a trajectory-location dual-tower architecture and MoE Transformer blocks to enhance prediction accuracy across diverse urban environments.
  • TrajGPT (Hsu et al., 2024): TrajGPT is a multitask transformer-based spatiotemporal generative model that frames controlled synthetic trajectory generation as a text infilling problem, ensuring spatiotemporal consistency in visit sequences.
  • TrajFM (Lin et al., 2024): TrajFM is a trajectory foundation model that emphasizes both region transferability and task transferability through its STRFormer backbone.
  • Unitraj (Zhu et al., 2024): UniTraj is a universal trajectory foundation model trained on large-scale, worldwide human mobility data, designed to be both task-adaptive and region-independent.
  • Difftraj (Zhu et al., 2023): DiffTraj is a spatiotemporal diffusion probabilistic model designed for GPS trajectory generation. It employs a Trajectory UNet (Traj-UNet) to reverse a noise process, reconstructing high-fidelity trajectories that preserve the original spatial distributions
  • TrajGAN (Goodfellow et al., 2014): TrajGAN introduces a GAN framework for trajectory generation, where trajectories perturbed with random noise are mapped into realistic sequences by a generator composed of linear and convolutional layers.
@@
  • Movesim (Feng et al., 2020): MoveSim is a generative adversarial network (GAN)-based framework for simulating human mobility patterns. It combines model-based and model-free approaches, utilizing SeqNet for sequential transition modeling and RegNet for incorporating urban structure effects, outperforming existing models in metrics like travel distance and location visits.
  • SAE (Malhotra et al., 2016): SAE is a standard RNN-based sequence-to-sequence architecture trained by minimizing the reconstruction error. The anomaly score for a sequence is determined according to how large its reconstruction error is.
  • GMVSAE (Liu et al., 2020): GMVSAE is a model designed for detecting anomalies in vessel trajectories by employing a Gaussian Mixture Variational Autoencoder, which analyzes historical data and established shipping lanes to identify vessels that deviate significantly from expected paths.

### v3

@@ 1 Introduction
Understanding human mobility patterns is a cornerstone for gaining insight into complex human behaviors and socio-economic dynamics (song2010modelling; gonzalez2008understanding). The vast digital traces from billions of personal devices have provided researchers with massive datasets of human mobility, capturing daily activities on an unprecedented scale (feng2018deepmove; yuan2022activity; xue2021mobtcast). The effort to find predictable patterns within this apparent randomness is not new; nearly two decades ago, studies revealed surprising regularities and high predictability in human mobility (song2010limits; schneider2013unravelling). This combination of massive data and intrinsic predictability creates the ideal conditions to move beyond task-specific statistical models and towards a more holistic and generalizable understanding of human movement (zhou2024urban).
Inspired by the transformative success of foundation models in language, a new wave of general-purpose models for mobility has emerged, such as UniST (yuan2024unist), UrbanDiT (yuan2024urbandit), TrajFM (lin2024trajfm), TrajGPT (hsu2024trajgpt), UniMob (long2025universal), GenMove (long2025one), and UniTraj (zhu2024unitraj). However, these pioneering efforts, while valuable, fall short of creating a true “GPT for mobility”. A detailed comparison, summarized in Table 1, reveals several practical limitations shared by current approaches. Specifically, they rely on inconsistent basic units to represent movement, from raw GPS coordinates to non-transferable Zone IDs, a fragmentation that directly hinders their geographic scalability and prevents cross-city pre-training. Furthermore, many of these models exhibit limited task flexibility and are constrained in their parameter scale, which prevents them from fully leveraging the rich patterns within large, diverse mobility datasets.
@@
TrajFM (lin2024trajfm) Scratch Raw GPS Irregular 18M
@@
TrajGPT (hsu2024trajgpt) Scratch Stay Point Irregular 20w
TrajGDM (chu2023trajgdm) Scratch Stay Point Irregular 28M
UniMove (han2025unimove) Scratch Stay Point Irregular 40M
GenMove (long2025one) Scratch Zone ID Regular 21M
UniMob (long2025universal) Scratch Zone ID Regular 6M
UniST (yuan2024unist) Scratch Zone ID Regular 30M
UrbanGPT (li2024urbangpt) LLMs Zone ID Regular -
UrbanDiT (yuan2024urbandit) Scratch Zone ID Regular 30M
@@
We argue these practical limitations stem from two fundamental yet unsolved barriers. The first barrier is the lack of a universal basic unit, which directly prevents the scaling of data across geographies. For instance, some models define tokens as aggregated spatio-temporal patches (e.g., UniST (yuan2024unist), UrbanDiT (yuan2024urbandit)), an approach too coarse to capture the semantics of individual movement. Others use raw GPS coordinates (e.g., UniTraj (zhu2024unitraj)), which are too fragmented to represent daily routines, or rely on city-specific Zone IDs (e.g., TrajGPT (hsu2024trajgpt), UniMob (long2024universal), GenMove (long2025one)), which are not universal and prevent the creation of large-scale, cross-city pre-training datasets. The second barrier is architectural. Human mobility, much like language, is a mixture of distinct behaviors shaped by varied intents and geographical contexts (han2025unimove). Existing models, with their ”one-fits-all” parameters, are forced to learn an inefficient, averaged representation of these diverse patterns, failing to capture the rich semantics of movement as data diversity grows. Recent LLM-based approaches (wang2023would; li2024more; feng2025agentmove; du2024trajagent; du2025cams) sidestep this by reasoning in a language space, a process that divorces the model from the data’s intrinsic spatio-temporal structure.
@@ 3.1 Unified location encoder
Existing approaches typically rely on discrete IDs to encode locations. However, such representations hinder cross-city transferability. To build a mobility foundation model, it is essential to develop a unified encoder that can generate consistent representations of locations across different cities. Existing research (hu2020universal; song2010universal; gonzalez2008understanding; lenormand2020importance) shows evidence that human mobility is influenced by individual intents, travelers tend to choose their next destination based on travel needs and spatial distance. To capture this behavior, we use the locations’ functional feature (POI distribution) and geographic feature (geographic coordinates) as core features for the encoder. In addition, mobility patterns are also influenced by broader social dynamics. To account for this, we incorporate the social popularity level of each location as another key feature. Formally, for all locations in a city, we construct three embeddings:
@@ 3.3 Next-step Output
Instead of direct classification, we model next-step prediction as a similarity search problem. Specifically, a mobility trajectory $S$ is encoded by the SAMoE Transformer to produce a trajectory representation $H_{\text{traj}}^{l}\in\mathbb{R}^{d}$ for a given length $l$, which acts as a query. Concurrently, we create a candidate database $\mathcal{D}\in\mathbb{R}^{N\times d}$ containing embeddings for all $N$ locations in the city, generated by a Deep & Cross Network (DCN) (wang2017deep) from their features ($E_{\text{poi}}$, $E_{\text{geo}}$, $E_{\text{pop}}$) as follows:
@@ B.1 Human Mobility Modeling
Human mobility modeling encompasses a broad spectrum of tasks, including mobility prediction, generation, and anomaly detection. These tasks jointly provide complementary perspectives for understanding, simulating, and evaluating human movement behaviors (song2010limits; schneider2013unravelling; fang2025unraveling). Early approaches were dominated by probabilistic models such as Markov chains and exploration-preferential return (EPR) models (gambs2012next; jiang2016timegeo), which captured mobility patterns from transition probabilities but struggled with the irregular and dynamic nature of human behavior. Recent advances in deep learning have significantly boosted the field, with recurrent neural networks (RNNs) (chen2020predicting; yuan2022activity), attention-based architectures (feng2018deepmove; yuan2023spatio), Transformers (corrias2023exploring; si2023trajbert), graph neural networks (GNNs) (sun2021periodicmove; terroso2022nation), and diffusion models (zhu2023difftraj; long2025one) being widely applied across different tasks. However, most of these models remain city-specific, requiring separate training for each dataset and thus limiting their generalizability. Inspired by the paradigm shift in natural language processing, a series of mobility foundation models have recently been proposed, including UniST (yuan2024unist), UrbanDiT (yuan2024urbandit), TrajFM (lin2024trajfm), TrajGPT (hsu2024trajgpt), UniMob (long2025universal), GenMove (long2025one), and UniTraj (zhu2024unitraj). These approaches attempt to leverage large-scale pre-training to improve few-shot and zero-shot performance, yet most still can not capture the universality and diversity of human mobility patterns across urban environments.
@@ B.2 Mixture of Experts
Mixture of Experts (MoE)  (shazeer2017outrageously; jordan1994hierarchical; jacobs1991adaptive) is a deep learning architecture that enhances model performance through a divide-and-conquer strategy. Its core idea lies in decomposing complex tasks and enabling collaborative processing by multiple specialized sub-models (experts), guided by a learnable gating network that dynamically allocates input data to the most relevant experts (zhou2022mixture). Each expert processes only a subset of inputs, significantly reducing computational costs, while the gating mechanism synthesizes final results through weighted aggregation of expert outputs (fedus2022switch; lepikhingshard). By maintaining the model’s parameter scale while improving inference efficiency, MoE has been widely adopted in large-scale language models (liu2024deepseek; du2022glam; lepikhingshard), enabling trillion-parameter model deployment via sparse activation. However, challenges such as expert load balancing and training stability persist in its implementation.
@@ D.1 Baselines
  • Markov (gambs2012next): The Markov model provides a statistical framework for characterizing how states evolve over time. It predicts future locations by estimating the transition probabilities between states.
  • LSTM (graves2012long): Long Short-Term Memory networks are highly effective for sequential modeling, as they capture long-term dependencies and temporal dynamics, making them particularly suitable for location prediction tasks.
  • Transformer (vaswani2017attention): The Transformer architecture leverages self-attention mechanisms to process sequential data efficiently, enabling the modeling of long-range dependencies in human mobility patterns.
  • DeepMove (feng2018deepmove): DeepMove is designed to address the challenges of human mobility prediction by jointly considering spatial, temporal, and personal preference factors that shape individual movement behaviors.
  • GetNext (yang2022getnext): GETNext integrates Graph Convolution Networks with Transformers to model global POI transition patterns, incorporating time-aware semantic embeddings for next-location recommendation.
  • CTLE (lin2021pre): The Context- and Time-aware Location Embedding model dynamically generates location embeddings by combining bidirectional Transformers with temporal encoding, capturing multi-faceted semantics of locations under varying spatial-temporal dynamics.
  • Unimove (han2025unimove): Unimove is a unified model for multi-city human mobility prediction, employing a trajectory-location dual-tower architecture and MoE Transformer blocks to enhance prediction accuracy across diverse urban environments.
  • TrajGPT (hsu2024trajgpt): TrajGPT is a multitask transformer-based spatiotemporal generative model that frames controlled synthetic trajectory generation as a text infilling problem, ensuring spatiotemporal consistency in visit sequences.
  • TrajFM (lin2024trajfm): TrajFM is a trajectory foundation model that emphasizes both region transferability and task transferability through its STRFormer backbone.
  • Unitraj (zhu2024unitraj): UniTraj is a universal trajectory foundation model trained on large-scale, worldwide human mobility data, designed to be both task-adaptive and region-independent.
  • Difftraj (zhu2023difftraj): DiffTraj is a spatiotemporal diffusion probabilistic model designed for GPS trajectory generation. It employs a Trajectory UNet (Traj-UNet) to reverse a noise process, reconstructing high-fidelity trajectories that preserve the original spatial distributions
  • TrajGAN (goodfellow2014generative): TrajGAN introduces a GAN framework for trajectory generation, where trajectories perturbed with random noise are mapped into realistic sequences by a generator composed of linear and convolutional layers.
@@
  • Movesim (feng2020learning): MoveSim is a generative adversarial network (GAN)-based framework for simulating human mobility patterns. It combines model-based and model-free approaches, utilizing SeqNet for sequential transition modeling and RegNet for incorporating urban structure effects, outperforming existing models in metrics like travel distance and location visits.
  • SAE (malhotra2016lstm): SAE is a standard RNN-based sequence-to-sequence architecture trained by minimizing the reconstruction error. The anomaly score for a sequence is determined according to how large its reconstruction error is.
  • GMVSAE (liu2020online): GMVSAE is a model designed for detecting anomalies in vessel trajectories by employing a Gaussian Mixture Variational Autoencoder, which analyzes historical data and established shipping lanes to identify vessels that deviate significantly from expected paths.